Carbon dioxide (CO2) is an abundant greenhouse gas
Answer #1

Lewis structure for Bicarbonate anion:

Note that you just missed the braket stating the -1 charge of the ion

electrons count is correct and perfectly placed

Q2

for

Nitrogen dioxide + water = nitric acid + nitrous acid

NO2 + H2O = HNO3 + HNO2

balance

2 NO2(g) + H2O(l) = HNO3(aq) + HNO2(aq)

Q3

Nitrogen dioxide gas into O2 and N2:

NO2(g) = N2(g) + O2(g)

balance

2NO2(g) = N2(g) + 2O2(g)